how many litres of 10% solution of acid should be added to 20 litres of a 60% solution of acid to obtain a 50% solution?
Answer by stanbon(75887)

how many litres of 10% solution of acid should be added to 20 litres of a 60% solution of acid to obtain a 50% solution?
-------------------
Equation:
acid + acid = acid
0.10x + 0.60*20 = 0.50(x+20)
------------------------
Multiply thru by 100 to get:
10x + 60*20 = 50x + 50*20
40x = 10*20
x = 5 litres (amt. of 10% solution needed)
=============
Cheers,
Stan H.